OpenStudy (anonymous):

is 2a-c/m the same as a-c/m +a

OpenStudy (anonymous):

Are the denominators the same for (a-c)/m and (a/1) ?

OpenStudy (anonymous):

We lets check it out \[ \frac{a-c}{m}+a\] We can only add fractions if they have the same base. So the common denominator would be m so lets multiply the top and bottom by m \[ \frac{a-c}{m}+a* \frac{m}{m}\] \[ \frac{a-c}{m}+\frac{a*m}{m}\] Now lets add the 2 fractions \[ \frac{a-c+am}{m}\]

OpenStudy (anonymous):

So is it the same?